Find the linear equation of the line that passes through (-3,2) and (5,-7).
Answer by stanbon(75887)

Find the linear equation of the line that passes through (-3,2) and (5,-7).
-------------
slope = (-7-2)/(5--3) = -9/8
--------------
Form is y = mx +b and you know y=-7 when x=5 and m=-9/8
----------------
-7 = (-9/8)*5 + b
-7 = -45/8 + b
-56/8 = -45/8 + b
b = -11/8
-----------------
EQUATION:
y = (-9/8)x -(11/8)
=========================
Cheers,
Stan H.